OpenStudy (anonymous):

If AB=7, what is the area between the two circles?

OpenStudy (anonymous):



OpenStudy (callisto):



OpenStudy (callisto):

Suppose the two circles are concentric. CB = 7/2 = 3.5 (line joining from centre perpendicular to chord bisects chord) By Pyth. Thm., \[r_2^2 = r_1^2 + 3.5^2\]So, \[r_2^2 - r_1^2 = 3.5^5\] Area between two circles = area of big circle - area of small circle \[= \pi (r_2^2 - r_1^2)\]The rest are substitution and calculations.

OpenStudy (callisto):

*facepalm* For the sixth line, it should be \[r_2^2 - r_1^2 = 3.5^2\]